# High Probability Bound for Cross-Learning Contextual Bandits with Unknown Context Distributions

### v1

@@ B.1 Decomposition
As we mentioned in Section 3.2, we decompose the regret as
@@
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\hat{\ell}_{t,c}\rangle}_{\textbf{ftrl}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\tilde{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{3}}$
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{4}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{5}}.$
@@ B.2 Upper Bounding $\operatorname{\mathbf{bias}}_{5}$
We first bound the fifth term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$. We decompose the fifth term into two components:
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\ell_{t,c}(\pi_{c})\right)$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
@@
    $=$ $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right].$
@@
    $\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=$ $\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}+\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=$ $\mathbb{E}\left[\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}\frac{\left(f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{1}{2}\gamma\right)}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]-\frac{\gamma}{f_{e}(\pi_{c})+\gamma}\mathbb{E}\left[F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
    $=\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}$
@@
    $=$ $\mathbb{E}\left[\left(\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq\sum_{c}\Pr(c)\mathbb{E}\left[\left(\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=\sum_{c}\Pr(c)(\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c})^{2}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq L\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq 4L\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{f_{e}(\pi_{c})+\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$ (Lemma 3)
    $\leq\sum_{c}\Pr(c)\frac{4L}{(f_{e}(\pi_{c})+\gamma)^{2}}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\mathbb{E}\left[\left(f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=\sum_{c}\Pr(c)\frac{4L}{(f_{e}(\pi_{c})+\gamma)^{2}}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\left(\mathbb{E}\left[\left(f_{e}(\pi_{c})-\hat{f}_{e}(\pi_{c})\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]+\frac{9}{4}\gamma^{2}\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
The second term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ has the following properties:
@@
    The sequence of random variables $\left\{\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$ forms a martingale difference sequence with respect to the tfiltration $\{\operatorname{\mathcal{H}}_{t}\}_{t}$.
@@
    Each random variable $\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ satisfies $\left|\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right|\leq\frac{1}{\gamma}$.
@@
    Each random variable $\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ satisfies
    $\mathrm{Var}\left[\sum_{c}\Pr(c)\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$ $\leq\sum_{c}\Pr(c)\mathrm{Var}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
    $=\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})-f^{2}_{e}(\pi_{c})}{\left(\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c})^{2}$
    $\leq\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})}{\left(\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c}).$
Applying Freedman’s inequality to the sequence of random variables $\left\{\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$ , we get that for each $\delta\in(0,1)$ and each $0<\lambda<\gamma$, with probability at least $1-\delta$, we have
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
$\leq$ $\lambda\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})}{\left(\hat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c})+\frac{1}{\lambda}\log(\frac{1}{\delta}).$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$

### v2

@@ B.1 Decomposition
As we mentioned in Section 4, we decompose the regret as
@@
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\widehat{\ell}_{t,c}\rangle}_{\textbf{ftrl}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\widetilde{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{3}}$
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{4}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{5}}.$
@@ B.2 Upper Bounding $\operatorname{\mathbf{bias}}_{5}$
We first bound the fifth term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle$. We decompose the fifth term into two components:
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\ell_{t,c}(\pi_{c})\right)$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
@@
    $=$ $\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right].$
@@
    $\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=$ $\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}-\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}+\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}-1\right)F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=$ $\mathbb{E}\left[\frac{f_{e}(\pi_{c})}{f_{e}(\pi_{c})+\gamma}\frac{\left(f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{1}{2}\gamma\right)}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]-\frac{\gamma}{f_{e}(\pi_{c})+\gamma}\mathbb{E}\left[F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
    $=\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}$
@@
    $=$ $\mathbb{E}\left[\left(\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq\sum_{c}\Pr(c)\mathbb{E}\left[\left(\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\frac{f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=\sum_{c}\Pr(c)(\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c})^{2}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq L\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $\leq 4L\sum_{c}\Pr(c)\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}\mathbb{E}\left[\left(\frac{f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma}{f_{e}(\pi_{c})+\gamma}F_{e}\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$ (Lemma 3)
    $\leq\sum_{c}\Pr(c)\frac{4L}{(f_{e}(\pi_{c})+\gamma)^{2}}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\mathbb{E}\left[\left(f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})-\frac{3}{2}\gamma\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
    $=\sum_{c}\Pr(c)\frac{4L}{(f_{e}(\pi_{c})+\gamma)^{2}}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\ell_{t,c}(\pi_{c})\left(\mathbb{E}\left[\left(f_{e}(\pi_{c})-\widehat{f}_{e}(\pi_{c})\right)^{2}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]+\frac{9}{4}\gamma^{2}\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
The second term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ has the following properties:
@@
    The sequence of random variables $\left\{\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$ forms a martingale difference sequence with respect to the tfiltration $\{\operatorname{\mathcal{H}}_{t}\}_{t}$.
@@
    Each random variable $\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ satisfies $\left|\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right|\leq\frac{1}{\gamma}$.
@@
    Each random variable $\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ satisfies
    $\mathrm{Var}\left[\sum_{c}\Pr(c)\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$ $\leq\sum_{c}\Pr(c)\mathrm{Var}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
    $=\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})-f^{2}_{e}(\pi_{c})}{\left(\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c})^{2}$
    $\leq\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})}{\left(\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c}).$
Applying Freedman’s inequality to the sequence of random variables $\left\{\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$ , we get that for each $\delta\in(0,1)$ and each $0<\lambda<\gamma$, with probability at least $1-\delta$, we have
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
$\leq$ $\lambda\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\frac{f_{e}(\pi_{c})}{\left(\widehat{f}_{e}(\pi_{c})+\frac{3}{2}\gamma\right)^{2}}l_{t,c}(\pi_{c})+\frac{1}{\lambda}\log(\frac{1}{\delta}).$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle$